OpenStudy (anonymous):

Evaluate the expression below when x = -4. 4x2 - 2x × 15

OpenStudy (alexwee123):

4(-4)^2 -2(-4)*15 =4(16)+8*15 =64+120 =184

OpenStudy (alexwee123):

questions? ^.^

OpenStudy (anonymous):

thanks! and you want moree?

OpenStudy (alexwee123):

more questions? o.0

OpenStudy (anonymous):

Evaluate the expression below when g = -6. -3g - 2 -(-4g + 10)

OpenStudy (alexwee123):

first sub in the -6 for g\[-3(-6)-2-(-4*(-6)+10) \]

OpenStudy (anonymous):

so justt solve that?

OpenStudy (alexwee123):

yep !

OpenStudy (alexwee123):

tell me what you get after you solve it

OpenStudy (anonymous):

30

OpenStudy (alexwee123):

nope try again :/

OpenStudy (alexwee123):

hint* you should get a negative number for the answer

OpenStudy (anonymous):

-14

OpenStudy (anonymous):

-18***?

OpenStudy (alexwee123):

ya you got it:-18 ^.^

OpenStudy (anonymous):

thx!

OpenStudy (alexwee123):

any more questions?

OpenStudy (anonymous):

not at the moment

OpenStudy (anonymous):

Find the value of y when x equals 1 in this equation. 2x + (-7y) = -54

OpenStudy (alexwee123):

@b.l.m hmm what do you think the first step is?

OpenStudy (anonymous):

solve it

OpenStudy (alexwee123):

yes but how would you solve it?

OpenStudy (anonymous):

puttin 1 into x?

OpenStudy (alexwee123):

yep then what would you do next?

OpenStudy (anonymous):

solve it

OpenStudy (alexwee123):

yep :) so can you do it for me?

OpenStudy (anonymous):

yupp

OpenStudy (alexwee123):

aight tell me what you get after you solve it

OpenStudy (anonymous):

-6?

OpenStudy (alexwee123):

close but y does not equal -6 \[2(1) + (-7y) = -54\]\[-7y=-56\]\[y=8\]

OpenStudy (anonymous):

thx

OpenStudy (alexwee123):

anytime ^.^